# Thermometer: Towards Universal Calibration for Large Language Models

### 20 Feb 2024 v1

@@ C.4 Implementation Details
In the implementation of thermometer, the input dimension of thermometer is set to be 2048 and 4096 for FLAN-T5-XL and LLaMA-2-Chat 7B , respectively. Correspondingly, the dimensions of the hidden layers in thermometer are set at 256 for FLAN-T5-XL and 512 for LLaMA-2-Chat 7B . To ensure that the output of thermometer remains positive, a Softplus activation function is adopted. The optimization of thermometer utilizes the AdamW optimizer, and all the hyper-parameter used for training is summarized in Table 15. The configuration is consistent across experiments conducted on MMLU, BIG-bench, and MRQA. All experiments used PyTorch. Experiments involving LLaMA-2-Chat 13B and LLaMA-2-Chat 70B were run on three A100 GPUs with 80GB memory each. All other experiments were run on a single Tesla V100 GPU with 32 GB memory.

### 27 Jun 2024 v2

@@ C.4 Implementation Details
In the implementation of thermometer, the input dimension of thermometer is set to be 2048 and 4096 for FLAN-T5-XL and LLaMA-2-Chat 7B , respectively. Correspondingly, the dimensions of the hidden layers in thermometer are set at 256 for FLAN-T5-XL and 512 for LLaMA-2-Chat 7B . To ensure that the output of thermometer remains positive, a Softplus activation function is adopted. The optimization of thermometer utilizes the AdamW optimizer, and all the hyper-parameter used for training is summarized in Table 16. The configuration is consistent across experiments conducted on MMLU, BIG-bench, and MRQA. All experiments are implemented in PyTorch using Tesla V100 GPU with 32 GB memory and Tesla A100 GPU with 40 GB memory.